OpenStudy (anonymous):

A system of equations is given below: y = –2x + 1 6x + 2y = 22 Which of the following steps could be used to solve by substitution? A. 6x + 2(–2x + 1) = 22 B. –2x + 1 = 6x + 2y C. 6(–2x + 1) + 2y = 22 D. 6(y = –2x + 1)

OpenStudy (anonymous):

plug in y

OpenStudy (anonymous):

so 6x+2(-2x+1)=22

OpenStudy (anonymous):

it would be A. y = 1/2 and x = 3.5

OpenStudy (anonymous):

thank you!